# Evaluating AI-Generated Essays with GRE Analytical Writing Assessment

### v2

@@ 3 Results
3.1 Human Scores
We recruited expert human raters from the GRE program to score the 200 essays (100 from each of the two writing prompts) generated by each LLM. The raters followed the same scoring framework as in the GRE scoring operation and provided a score ranging from 0 to 6 for each essay (see Appendix for the definition of the scores). The source of the essay, i.e., written by humans or AI, has been deliberately hidden from the human raters, so they were unaware that some essays were generated by AI. The score distribution is show in Table 2. We observed that essays generated by the proprietary LLMs (e.g., GPT models and Gemini) generally received higher scores from human raters. Newer models in 2024 generally do better than the LLMs in 2023. The average scores are between 4 and 5, which correspond to ”presents a competent analysis of the issue and conveys meaning with acceptable clarity” and ”generally thoughtful, well-developed analysis of the issue and conveys meaning clearly” respectively according to the GRE scoring framework.
3.2 e-rater® Scores
In addition to human scores, we applied the e-rater® engine to score each AI-generated essay. The score distribution is listed in Table 2.
@@ 3.1 Human Scores
$faLock \faUnlock \faLock \faLock \faLock \faUnlock \faUnlock \faUnlock \faUnlock$
@@
[…]
Compared with human scores, e-rater® tended to give higher scores to AI-generated essays, though the proprietary LLMs still outperform the open-source counterparts. This is unsurprising since e-rater® was developed and calibrated based on human-written essays and showed good agreement with human scores J. Chen et al. (2016). The current language features, such as grammar, usability, mechanics, and so on, are sufficient for scoring human-written essays but fall short of properly scoring AI-generated essays, which display almost impeccable language usage but with shallow meanings. In contrast, human raters adeptly recognize these shortcomings and assign more accurate scores to AI-generated essays.
3.3 Essay Length, Similarity, Language Features, and Perplexity
Figure 1: Distribution of the length of the generated essays. The red vertical dashed line indicates the mean essay length.
Essay Length Adherence
We introduced a length requirement, i.e., about 500 words, when prompting LLMs to generate essays. However, not all LLMs follow the instructions strictly. Figure 1 shows the distribution of the lengths of the generated essays in terms of the number of words. Overall, we observe that newer models exhibited smaller variances and better alignment with the 500-word instruction. For more recent models in 2024, GPT-4o and Gemini generated essays with lengths slightly greater than 500, while Llama3-8b generated an average of 460 words. For older models, close-sourced models performed better than open-sourced models (Mistral, Vicuna, Koala, and Llama) in following the length instruction. Essays generated by Llama displayed the largest length variability, indicating that the model struggled to consistently produce long and high-quality essays, which echoed the findings on essay quality.
Essay Similarity
The pairwise similarity of essays generated by the same LLM under the same prompt reflects consistency and variability under identical generation conditions (e.g., prompt, temperature). There are two commonly known types of similarities between texts: verbatim similarity, which describes how closely two texts match in exact wording or surface structure, and semantic similarity, which shows how similar two texts are in meaning, regardless of specific word choices. Comparing these two types of similarities for pairwise essays generated by LLMs under the same prompt could reveal how well the LLMs maintain meaning across different expressions of the same content.
For semantic similarity, we first convert each essay into a document vector using a popular neural embedding model jina-embedding-v3 model (Sturua et al., 2024). This embedding model was chosen based on its high performance on text similarity benchmarks of the MTEB English leaderboard [^1] https://huggingface.co/spaces/mteb/leaderboard and its superior performance on handling long documents. Once the essays are converted to vectors, we compute the cosine similarity between pairs of vectors corresponding to essays generated from the same LLM under the same prompt. Figure 2 shows the distribution of cosine similarity from two prompts. We observed that essays generated by the Llama model displayed the least similarity, comparable to human-written essays. Models such as GPT-4, Gemini, and GPT-4o tend to have distributions with higher mean cosine similarity, indicating that their generated essays are more semantically similar to each other, a sign of good alignment to the writing prompts. The distributions between the two prompts differ slightly, with some models demonstrating higher density and peaks for one prompt than another (i.e., GPT-4o has a higher density for prompt 1, while Llama3-8b obtains the highest mean similarity on prompt 2).
For verbatim similarity, we represent each essay as a trigram vector, a method proven effective in capturing textual similarities for test security applications (Choi et al., 2024). Based on the trigram vectors, we computed pairwise cosine similarities for each pair of essays from the same LLM and prompt. Figure 3 shows the distribution of the similarity. Overall, human essays displayed the lowest similarity, while different LLMs showed different levels of similar texts. Despite the essays from the same LLM and prompt showing a high level of semantic similarity, the verbatim similarity is much lower, though they are generally higher than that from human-written essays.
Figure 2: Semantic similarity between pairs of essays generated by the same LLM and prompt, measured using the cosine similarity of their embedding vectors.
Figure 3: Verbatim similarity between pairs of essays generated by the same LLM and prompt, measured using the cosine similarity of their trigram vector representation.
Language Features
The language features from e-rater® allow us to compare more detailed aspects of the generated essays across the LLMs. Table 3 shows the mean and its standard error of the language features. It’s important to note that grammar, mechanics, usage, and style features are negative by convention, as they represent errors—the closer the value is to 0, the fewer the errors. In contrast, organization, development, and word complexity features are positive, with higher values indicating better performance in these areas. To visualize the differences across LLM-generated and human-written essays, we apply MinMax scaling to bring all features into the same range (0 to 1) and show them in Figure 4. Based on the results, we have the following observations:
Humans are doing better in development but show weaker performance in grammar, mechanics, and usage. This suggests human writing is strong in structure and coherence but prone to technical errors compared to LLMs. The GPT family (particularly GPT-4o and GPT-4) shows strong performance in grammar, mechanics, and usage, with wide coverage on the radar plots, indicating an all-around high performance with fewer errors and more fluent text generation. The 2024 models generally perform better than the 2023 ones, suggesting steady progress in these LLMs.
Figure 4: Comparison of the language features (after MinMax scaling) of the essays from different LLMs and humans. The ones in blue colors are models in 2024 and the ones in green colors are those from 2023. The bigger the value, the better the essay quality in that dimension.
[…]
mid-2024
mid-2023
Perplexity Distribution
Perplexity and perplexity-based features have been used to detect AI-generated essays. Specifically, it has been shown that essay perplexity alone can distinguish between ChatGPT-generated and human-written essays in high-stakes writing assessment (Hao & Fauss, 2024). In this study, we used GPT-2 as the underlying language model and adopted the implementation of perplexity by HuggingFace. [^2] https://huggingface.co/docs/transformers/perplexity The results are shown in Figure 5. The perplexity of human-written essays is much higher than that of AI-generated essays. Chronically, the newer models (GPT-4o, Gemini) showed slightly higher perplexity than older models, which suggests the essays generated by these models deviated further from the GPT-2 baseline model than the other LLMs. The perplexity of essays from Llama shows the largest variability, indicating that the essays may not be consistently generated.
Figure 5: Boxplot of the essay perplexity distribution for different models. The ones in blue colors are models in 2024 and the ones in green colors are those from 2023. The bigger the value, the better the essay quality in that dimension.
4 Detectability of AI-generated Essays
As the quality of essays generated by LLMs continues to improve, concerns are rising in education about students submitting AI-generated assignments, posing a significant challenge to academic integrity. The detection of AI-generated essays in large-scale assessments has been a widely discussed topic (Hao & Fauss, 2024; Jiang et al., 2024a; Yan et al., 2023), with most previous work focusing on detecting essays from a single LLM. However, with new and more powerful LLMs emerging every few months, it becomes increasingly difficult to identify which LLMs are being used. This highlights the importance of examining the detectability of AI-generated essays across multiple LLMs.
In particular, we consider two scenarios: within-model detectability and cross-model detectability. In the within-model scenario, we assess how well classifiers trained on essays generated by a specific LLM can detect essays produced by that same LLM. In the cross-model scenario, we examine how effectively a classifier trained on data from one LLM can detect essays generated by other LLMs. In both cases, the classifiers were developed using the seven language features in Table 3 and perplexity-based features (including the whole essay perplexity score, the mean, median, max, min, and 10, 20, 30, …, 90 percentiles of the sentence perplexity in each essay) respectively. We experimented with multiple popular classifiers, including XGBoost (T. Chen & Guestrin, 2016), Random Forest (Breiman, 2001), Gradient-Boosted Machine (Friedman, 2001), and Support Vector Machine (Cortes & Vapnik, 1995), and found that XGBoost consistently gave the best results. So, we only show the results of the XGBoost classifier.
@@ 4 Detectability of AI-generated Essays
For the within-model scenario, the results in Table 4 show that classifiers using perplexity features slightly outperformed those using the e-rater® language features, although both performed well overall. Notable exceptions include the GPT-4o and Gemini models, where language features are slightly more effective. This suggests that essays generated by these typical LLMs can be reliably detected in the context of GRE analytical writing as long as the LLM used is known. It’s important to note that these detection results apply to AI-generated essays without human revisions. If AI-generated text is revised by humans or mixed with human-written content, detection accuracy is expected to decrease Hao & Fauss (2024); Hao et al. (2024).
@@
Table 4: Performance of the XGBoost classifier with different feature sets: e-rater® language features, perplexity feature on the binary classification task. Results are based on the average of five-fold cross-validation.
The results for the cross-model case are shown in Figure 6. Overall, perplexity feature-based classifiers performed slightly better than language feature-based models when the training and testing data came from LLMs of the same year. They also excelled when trained on data from 2024 LLMs and tested on 2023 LLM-generated data. On the other hand, language feature-based classifiers did much better when they were trained on data from 2023 LLMs and tested on data from 2024 LLMs. Surprisingly, classifiers trained on GPT-4o data using perplexity-based features achieved very high accuracy in detecting essays from most LLMs, with the exception of essays generated by Gemini and the 2023 version of Llama. This indicates that perplexity features are generally effective across models but may struggle with certain newer or more distinct language patterns in specific LLMs.
(a) Language feature-based classifiers
(b) Perplexity feature-based classifiers
Figure 6: Detection accuracy when training the detector with one model and evaluating it on essays generated by different models. The left panel presents results using language-based features in the detector, while the right panel shows results using perplexity-based features. Results are based on the average of five-fold cross-validation.
@@ 5 Discussion
In this paper, we conducted a large-scale empirical evaluation of ten top LLMs by using GRE analytical writing assessment. Our goal is to establish benchmarks for the quality of essays generated by AI in cases that require high-level analytical thinking skills. Our results indicate that proprietary LLMs, such as GPTs and Gemini, consistently produced higher-scoring essays compared to open-source models, aligning with existing evaluations that show proprietary models outperforming their open-source counterparts. Additionally, we observed a notable improvement in essay scores from 2024 models compared to 2023 models, reflecting the continuous efforts within the AI community to create increasingly capable models for generating higher-quality text.
Further analysis of the language features computed by the e-rater® engine revealed that AI-generated essays better-controlled grammar, usage, and mechanics errors than human-written essay baseline. All AI-generated essays show a lower perplexity than the baseline of human-written essays. These findings provide information for improving future AI text generation, emphasizing crafting more meaningful texts alongside its already high-quality linguistic characteristics.
@@
Additionally, we have devised several machine learning classifiers, using language features from e-rater® and perplexity features to distinguish between human-written and AI-generated essays. The performance of all these classifiers is promising, suggesting that AI-generated essays responding to GRE analytical writing prompts can be reliably identified. However, these results apply mainly when no further revisions have been made to the AI-generated essays. Nevertheless, these findings establish a benchmark for the detection accuracy of such essays within the context of GRE analytical writing, and the findings are consistent with findings from other writing assessments (Yan et al., 2023; Hao & Fauss, 2024).
@@
To summarize, this study has evaluated the essays generated by LLMs using the GRE analytical writing framework and benchmarked various properties under controlled conditions. We hope the findings will help the assessment community understand the status of AI-generated essays and the potential implications of writing assessments. Meanwhile, we hope the findings can also help the LLM community think about ways to enhance the quality of model outputs through improved training and fine-tuning techniques.
@@
The current study has several limitations. Firstly, LLMs are constantly evolving, and it is challenging to enumerate all available advanced models. Whether and how the findings of this paper will change as LLMs change needs follow-up investigation, which we will report in future work.
@@
Lastly, the variability of the generated essay is controlled by hyper-parameters, such as the temperature of the LLMs. Different LLMs may have different baseline settings for these hyper-parameters, which leads to varying degrees of variability in generated essays even when using the same hyper-parameter values. Therefore, a systematic investigation of the quality of essays for different hyper-parameters is imperative, and we will report the findings in future work.
Acknowledgment
YZ thank the ETS Research Summer Internship Program to fund the project. We also thank ETS Research Institute for funding the project.

### v4

@@ 4 Results
4.1 Comparing Characteristics of Essays
Figure 1: Distribution of the length of essays generated by different LLMs.
Essay Length Adherence. We introduced a length requirement, i.e., about 500 words, when prompting LLMs to generate essays. However, not all LLMs follow the instructions strictly. Figure 1 shows the distribution of the lengths of the generated essays in terms of the number of words. Overall, we observe that newer models exhibited smaller variances and better alignment with the 500-word instruction. For more recent models in 2024, GPT-4o and Gemini generated essays with lengths slightly greater than 500, while Llama3-8b generated an average of 460 words. For older models, close-sourced models performed better than open-sourced models (Mistral, Vicuna, Koala, and Llama) in following the length instruction. Essays generated by Llama displayed the largest length variability, indicating that the model struggled to consistently produce long and high-quality essays, which echoed the findings on essay quality.
Essay Similarity. For the pair-wise semantic similarity of essays generated by different LLMs under the same prompt, the results are shown in Figure 2.
Figure 2: Semantic similarity between pairs of essays generated by the same LLM and prompt.
We observed that essays generated by the Llama model displayed the least similarity, comparable to human-written essays. Models such as GPT-4, Gemini, and GPT-4o tend to have distributions with higher mean cosine similarity, indicating that their generated essays are more semantically similar to each other. The distributions between the two prompts differ slightly, with some models demonstrating higher density and peaks for one prompt than another (i.e., GPT-4o has a higher density for prompt 1, while Llama3-8b obtains the highest mean similarity for prompt 2). For the verbatim similarity, the results are shown in Figure 3.
Figure 3: Verbatim similarity between pairs of essays generated by the same LLM and prompt.
Overall, human essays displayed the lowest similarity, while different LLMs showed varying levels of similarity in their generated texts. Despite the essays from the same LLM and prompt showing a high level of semantic similarity, the verbatim similarity is much lower, though they are generally higher than that from human-written essays.
[…]
mid-2024
mid-2023
Language Features. The language features from e-rater® allow us to compare more detailed aspects of the generated essays across the LLMs. Table 2 shows the mean and standard error of the language features. It is important to note that grammar, mechanics, usage, and style features are negative by convention, as they represent errors: the closer the value is to 0, the fewer the errors. In contrast, organization, development, and word complexity features are positive, with higher values indicating better performance in these areas. To visualize the differences across AI-generated and human-written essays, we apply MinMax scaling to normalize all features to the same range (from 0 to 1) and present them in Figure 4. Based on the results, we have the following observations.
Figure 4: Comparison of the language features of the essays from different LLMs and humans. The ones in blue colors are models in 2024 and the ones in green colors are those from 2023. The original values of each feature from multiple LLMs have been rescaled using MinMax scaling, which transforms each value to a range between 0 and 1 based on the minimum and maximum observed values by $x^{\prime}=\frac{x-\min(x)}{\max(x)-\min(x)}$. After the scaling, higher values indicate better essay quality in that dimension.
Humans demonstrate stronger performance in development but show weaker performance in grammar, mechanics, and usage. This suggests that human writing is strong in terms of structure and coherence but prone to technical errors compared to LLMs. The GPT family (particularly GPT-4o and GPT-4) shows strong performance in grammar, mechanics, and usage, with wide coverage on the radar plots, indicating an all-around high performance with fewer errors and more fluent text generation. The 2024 models generally perform better than the 2023 models, suggesting steady progress in these LLMs.
Figure 5: Boxplot of the essay perplexity distribution for different models. The ones in blue colors are from essays generated by 2024 LLMs and the ones in green colors are those from 2023 LLMs.
Perplexity. Figure 5 compares the distribution of perplexity computed from essays generated by LLMs and written by humans. The perplexity of human-written essays is much higher than that of AI-generated essays, meaning human writing is less predictable to language models. Humans often use unexpected phrasing, varied structures, and idiosyncratic styles, while AI text tends to follow smoother, more statistically likely patterns. Higher perplexity thus reflects the greater variability and originality of human language compared with machine-generated text. Chronically, the newer models (GPT-4o, Gemini) showed slightly higher perplexity than older models, which suggests that the essays generated by these models deviated further from the GPT-2 baseline model than the other LLMs. The perplexity of essays from Llama shows the largest variability, indicating that the essays may not be consistently generated. These perplexity values serve as useful reference points for identifying essays generated by different LLMs.
4.2 Comparing Scores
@@
b b b b b
@@
Table 3: The percentage of essays in different score levels based on human raters and e-rater®. The symbols b and indicate open source and proprietary LLMs, respectively. The performance of the best and second-best models, as measured by the average of different scoring systems, is indicated in bold and underlined, respectively.
Table 3 presents the score distribution. For scores assigned by human raters, essays generated by proprietary LLMs, such as GPT models and Gemini, generally received higher scores than those produced by open-source models. Additionally, essays from newer models released in 2024 tended to receive higher scores than those from models introduced in 2023. The average scores of essays from top LLMs are between 4 and 5, which correspond to “presents a competent analysis of the issue and conveys meaning with acceptable clarity” and “generally thoughtful, well-developed analysis of the issue and conveys meaning clearly,” respectively, according to the GRE scoring framework. Compared to human raters, the e-rater® engine consistently assigned higher scores to AI-generated essays, and the essays from proprietary models also get higher e-rater® scores compared to those from open-source models. The inflated e-rater® score reveals important limitations in using the current system to evaluate essays generated (or significantly assisted) by AI.
As noted earlier, the e-rater score is determined by two parts: the language features and their weights, which are calibrated based on human-written essays. Since LLMs generate text based on learned probabilities from large language corpora, they tend to perform systematically better on many language features, such as grammar, mechanics, and usage. When e-rater® continues to apply the same feature weights originally optimized for human-written essays, it produces systematically higher scores for AI-generated content. This is not necessarily a flaw, given that it simply reflects the fact that AI-generated essays are doing much better in these aspects.
On the other hand, the language features used by e-rater® have traditionally served as reliable indicators of writing quality in human-authored essays, as strong language use often correlates with coherent reasoning and meaningful content. However, this relationship weakens with AI-generated texts, which appear fluent and with fewer errors in grammar, mechanics, and usage but often lack substance and depth. As a result, the existing set of language features used by e-rater does not cover deeper aspects of writing, such as logical and analytical reasoning, which are closely related to specific contexts. Without features that reflect these qualities, e-rater® is unable to evaluate the full depth of AI-generated writing. In contrast, experienced human raters can recognize the issues in AI-generated essays and adjust their scores accordingly. This mismatch in evaluative capability contributes to the observed discrepancy in scores between human raters and e-rater.
This discrepancy in scores between human raters and e-rater underscores the need for and possible directions to improve e-rater (or similar automated scoring engines) to better handle essays in the AI era, where AI-generated and AI-assisted writing is increasingly common. The improvements can be made in two areas: developing new features that cover deeper levels of thinking in the essay and recalibrating the feature weights to contribute to the final score. We want to emphasize that this does not necessarily mean the current version of e-rater® is flawed for its intended use for scoring human-written essays. It remains valid and reliable in that setting, where the existing language features are generally well aligned with writing quality. Its limitations arise only when the system is applied to texts created or influenced by LLMs, a scenario likely to become increasingly prevalent in the near future.
4.3 Comparing the Detection of AI-generated Essays
@@
To address RQ3, we compare the performance of the detection of AI-generated essays in two different scenarios. For the within-model scenario, where the detectors are trained and tested using essays generated from the same LLM, Table 4 shows that, overall, the classifiers based on perplexity features slightly outperformed those using the e-rater language features. Notable exceptions include the GPT-4o and Gemini models, where language features are slightly more effective. These findings suggest that when the LLM is known, essays generated by that model can be accurately detected in the context of GRE analytical writing. It is important to note that these detection benchmarks reflect only the detection of AI-generated essays without human revisions. If AI-generated text is revised by humans or mixed with human-written content, detection accuracy is expected to decrease (Hao & Fauss, 2024; Hao, von Davier, et al., 2024).
@@
Table 4: Performance of the XGBoost classifier with different feature sets: e-rater® language features, perplexity feature on the binary classification task. Note: Results are based on the average of five-fold cross-validation.
For the cross-model scenario, where the classifier is trained using essays generated by one LLM and tested on essays generated by a different LLM, Figure 6 and 7 how the results based on five-fold cross-validation. Overall, perplexity feature-based classifiers performed slightly better than language feature-based models when the training and testing data came from LLMs of the same year. They also excelled when trained on data from 2024 LLMs and tested on 2023 AI-generated data. On the other hand, language feature-based classifiers performed significantly better when trained on data from 2023 LLMs and tested on data from 2024 LLMs. Surprisingly, classifiers trained on GPT-4o data using perplexity-based features achieved very high accuracy in detecting essays from most LLMs, with the exception of essays generated by Gemini and the 2023 version of Llama. This suggests that perplexity features are generally effective across models but may struggle with certain newer or more distinct language patterns in specific LLMs.
Figure 6: Detection accuracy for classifiers trained on e-rater® language features
Figure 7: Detection accuracy for classifiers trained on perplexity-based features
@@ 5 Discussion
Over the past three years, especially since the release of ChatGPT, LLMs and generative AI have advanced at an unprecedented pace. Increasingly powerful LLMs are now widely used in various writing tasks and have been integrated into commonly used tools such as Microsoft Word and Google Docs. This widespread adoption signals the beginning of an era of co-created writing, where human and AI contributions are increasingly intertwined. This shift raises fundamental questions for writing assessment, including what aspect of writing should be assessed, how it should be assessed, and how academic integrity can be maintained in an AI-mediated writing environment.
@@
In this paper, we conducted a large-scale empirical study of AI-generated essays through the lens of the GRE Analytical Writing assessment to deepen understanding of their characteristics and examine their implications for two critical components of writing assessment: automated scoring and academic integrity. We compared the characteristics of essays generated by 10 popular LLMs and humans and observed that proprietary LLMs, such as GPTs and Gemini, consistently produced higher-scoring essays compared to open-source models, aligning with existing evaluations that have shown proprietary models outperforming their open-source counterparts. Additionally, we observed a notable improvement in essay scores from 2024 models compared to 2023 models, reflecting the continuous efforts within the AI community to create increasingly capable models for generating higher-quality text. Further analysis of the language features computed by the e-rater® engine revealed that AI-generated essays better controlled grammar, usage, and mechanics errors than human-written essay baseline. All AI-generated essays exhibit a lower perplexity than the baseline of human-written essays, which substantiates that perplexity is a crucial indicator for detecting AI-generated essays across LLMs. These findings provide valuable insights for improving future AI text generation, emphasizing the crafting of more meaningful texts alongside their already high-quality linguistic characteristics.
Another notable observation is that e-rater® systematically assigns higher scores to AI-generated essays compared to human raters. This discrepancy highlights limitations in current automated scoring systems when applied to AI-generated or AI-assisted writing. It also suggests directions for future improvements, such as developing new features that capture deeper meaning and reasoning and recalibrating the weighting of existing language features to more accurately reflect their contribution when AI tools can effectively handle them.
@@
Additionally, we examined the accuracy of machine learning classifiers that use language features from e-rater® and perplexity-based features to distinguish between human-written and AI-generated essays, considering both within-model and cross-model scenarios. The classifiers demonstrated strong performance, indicating that essays generated by AI in response to GRE Analytical Writing prompts can be reliably identified. It is important to note, however, that these results apply specifically to essays that are entirely generated by AI. When human edits are introduced or when the writing involves more iterative human-AI collaboration, detection becomes more challenging. Nonetheless, these findings provide a useful benchmark for detection accuracy in the context of GRE Analytical Writing, and the results are consistent with findings from other writing assessments (Hao & Fauss, 2024; Yan et al., 2023) .
@@
In summary, this study empirically examined essays generated by several major LLMs in the context of GRE Analytical Writing, benchmarked their characteristics, and analyzed the implications of AI-generated or assisted essays for automated scoring and academic integrity in the AI era. These findings aim to inform the assessment community about the current capabilities of AI-generated writing and its impact on writing evaluation practices. In the meantime, we hope the findings can also help the LLM community think about ways to enhance the quality of model outputs through improved training and fine-tuning techniques.
@@
This study has several limitations. First, LLMs are rapidly evolving, making it challenging to capture the full range of advanced models available at any given time. As a result, the characteristics and quality of AI-generated essays are subject to continuous change. Ongoing research will be necessary to monitor how the findings presented here may shift as newer models are developed and adopted.
@@
Third, the variability of the generated essay is controlled by hyper-parameters, such as the temperature of the LLMs. Different LLMs may have different baseline settings for these hyper-parameters, which leads to varying degrees of variability in generated essays, even when using the same hyperparameter values. Therefore, a systematic investigation of the quality of essays for different hyper-parameters is imperative, and we will report the findings in future work.
Finally, we proposed several high-level strategies for improving automated scoring systems when evaluating AI-generated or AI-assisted essays. However, we did not provide detailed specifications for new features or weight adjustment methods. This is partly because the quality of AI-generated essays is highly sensitive to the specific language model and prompting strategy employed. Additionally, our suggestions are based on a specific scoring engine, e-rater®, which may not fully reflect the challenges faced by other automated scoring systems. These issues warrant careful follow-up, and we plan to report further findings in future work.